What's the standard form of the equation y+4=-12/7(x-1)
Answer by unlockmath(1688)

Hello,
First, let's distribute y+4=-12/7(x-1)like this:
y+4=-12/7x+12/7 Now subtract 4 from both sides to result in:
y=-12/7x -16/7 This is standard form of y=mx+b
RJ
www.math-unlock.com